OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#7651 — lil-3/4-n5
Scheduled Maintenance Report for Web Cloud
Completed
We still have ARP issues on the switches managing the ADSL. Cisco has fixed the problem with a new image compiled especially for OVH.

We are going to apply it on the switches in Lille in order to check whether the problem is fixed.



Update(s):

Date: 2012-11-20 06:21:49 UTC
Done.

Date: 2012-11-20 06:21:40 UTC
Done. We are updating lil-3-n5.

Date: 2012-11-20 06:21:19 UTC
We are dealing with lil-4-n5. We are switching the traffic on lil-3-n5.
Posted Nov 20, 2012 - 06:20 UTC